OpenStudy (anonymous):

One of the offices at the company was being repainted. Riley’s father challenged him to figure out how many quarts of paint would be needed. The dimensions of the office are 16 ft by 12 ft by 10 ft. The office floor, ceiling, and door would not be painted and there are no windows in this office. The door opening measures 3 ft by 8 ft. One quart of paint covers 100 ft2. How many quarts of paint would be needed to paint the room, if only full quarts can be purchased?

OpenStudy (anonymous):

@KamiBug last one thanks for helping me so much thanks so much:))))

OpenStudy (anonymous):

(wall area - door area)/100 qts

OpenStudy (anonymous):

im not sure i can do this without you

OpenStudy (kamibug):

I think the total area of the four walls is 700ft^2. But, I'm not sure b/c the question doesn't specify which of the dimensions are height, width, or length... The area of the door is 3*8=24 Subtract that from the total wall area: 700-24=676 Divide that by 100 to find how many quarts you'll need. 676 / 100 = 7 quarts But, again, this might be wrong because I'm not sure I calculated the total wall area correctly, sorry. :[